OpenStudy (anonymous):

Graph the vertex:

OpenStudy (anonymous):

\[y=-x^2-2x\]

OpenStudy (anonymous):

I want to guess and say it's about.... (-1,-1) or so?

OpenStudy (anonymous):

first coordinate of the vertex is \[-\frac{b}{2a}\] in your case it is \[-\frac{-2}{2\times (-1)}=-1\]

OpenStudy (anonymous):

replace x by -1 get \[y=-(-1)^2-2\times (-1)=-1+2=1\]

OpenStudy (anonymous):

(1,-1)

OpenStudy (anonymous):

Why wouldn't the y be +1 and the x be -1?

OpenStudy (swag):

1

OpenStudy (anonymous):

Yea... I think it's just (1,-1)

OpenStudy (anonymous):

Er (-1,1)*

OpenStudy (anonymous):

(1,-1)

OpenStudy (anonymous):

Explain yourself please. :I

OpenStudy (anonymous):

gt sme numbers and apply them to the equation and see..

OpenStudy (anonymous):

*shurgs* I did... I got (-1,1)...

OpenStudy (anonymous):

yeah (1-,1)